OpenStudy (anonymous):

What is the y-intercept of the line? y=-7x+1 y-intercept =

Directrix (directrix):

The y-intercept of the line is where the line crosses the y-axis. To get on the y-axis, the x-coordinate of the coordinates of the y-intercept has to be 0.

Directrix (directrix):

So, take y=-7x+1. In place of x, put 0. Solve y = - 7 * (0) + 1 for y to get the y intercept. Post what you get and we can compare answers, okay?

Directrix (directrix):

What do you get when you multiply -7 times 0 ? @Christiana19199

OpenStudy (anonymous):

I get 0

OpenStudy (anonymous):

0+1 is 1

Directrix (directrix):

@Christiana19199 Look at this: y = - 7 * (0) + 1 ---> 7*0 = 0 So, y = - 7 * (0) + 1 = 0 + 1 = 1.

Directrix (directrix):

The y-intercept is 1. In coordinates, it is (0,1). Click on the attachment to see the graph.

Directrix (directrix):

@Christiana19199 Do you have a question about this?